OpenStudy (anonymous):

What is the equation of the line that contains the points with (x,y) coordinates (-3,7) and (5,-1)?

OpenStudy (anonymous):

The equation of a line in slope-intercept form is: y = mx + b You need to find two things, m and b m is the slope, and the slope is the difference in y over the difference in x Take your two given points (-3, 7) and (5, -1) the difference in y is (-1) - 7 = -8 the difference in x is 5 - (-3) = 8 m = (-8)/8 = -1 Now insert one point and m into the equation. y = mx + b Let's use (5, -1), and m = -1 -1 = -1(5) + b -1 = -5 + b Add 5 to both sides 4 = b Since we now know m = -1 and b = 4, we insert those values into the equation of a line: y = mx + b y = -x +4 Final answer

OpenStudy (anonymous):

Thank you very much, I really appreciate it